OpenStudy (anonymous):

express each complex number in polar form 5.2i another problem 6+4i

OpenStudy (anonymous):

is 5.2i 5 + 2i or 5 - 2i

OpenStudy (anonymous):

nope 5.2i

OpenStudy (anonymous):

A complex number z in the form x+y*i is described the same way any other polar equation is:\[x = R \cos\theta, y = R \sin\theta\] Where \[R = |z| = \sqrt{x^2 + y^2}.\] So, in this case, x = 0, and y = 5.2. So, R is going to be 5.2. When you want to express it in polar coordinates, you end up with the form \[z = R (\cos\theta + i*\sin\theta)\] (because the y component is multiplied by i).